# Galaxy Morphologies at Cosmic Noon with JWST: A Foundation for Exploring Gas Transport with Bars and Spiral Arms

### v1

@@ 4 Morphological evolution
In this section, we present and interpret the results of our visual classification of morphologies, with a discussion on the various found trends across redshift.
(a) Spiral fractions defined as $f_{\mathrm{spiral}}=N_{\mathrm{spiral}}/N_{\mathrm{total}}$ for comparison with other studies. Green dots and triangles represent the JWST-based results from (Kuhn et al., 2024), where the white circles show the directly measured spiral fractions, and the filled triangles indicate the values after applying corrections for observational effects. Orange dots represent the fractions inferred from HST imaging in the CANDELS fields by (Margalef-Bentabol et al., 2022).
(b) Bar fractions ($f_{\mathrm{bar}}=N_{\mathrm{bar}}/N_{\mathrm{disk}}$) from the subsample of disks with $e<0.5$ where the isophotal fitting analysis is reliable. Orange dots show the results from Guo et al. (2024), while green dots correspond to the fractions inferred by Le Conte et al. (2024), with green triangles indicating the fractions inferred using only their sample of strongly barred galaxies. Grey triangles represent simulated bar fractions from Liang et al. (2023) for a set of JWST filters, given an intrinsic $f_{\mathrm{bar}}$ of 68% at $z\approx 0$.
Figure 10: Comparison of spiral (left) and bar (right) fractions as a function of redshift. Blue markers represent results from our visual classifications, with $3\sigma$ error bars, which, similarly to the disk fraction errors, are estimated from a binomial distribution $\sigma=\sqrt{p(1-p)/N}$, where $p$ is the fraction and $N$ is the total number of galaxies in the parent sample.
@@ 4.1 Disks
Figure 9 compares the disk fractions $f_{\mathrm{disk}}=N_{\mathrm{disk}}/N_{\mathrm{total}}$ derived from our study, with the kinematic analysis of KMOS3D (Wisnioski et al., 2015) and a morphological classification by Ferreira et al. (2023). We find that our large disk fractions (90% and 72% in the low- and high-redshift bins, respectively) are consistent with the disk fractions estimated from the kinematic analysis of (Wisnioski et al., 2015) across the same redshift range. Additionally, our low-redshift results are also consistent with the findings of Tiley et al. (2021) (KGES) for their sample of 288 galaxies at $1.81.2<z<$ 1.8, where the $z\approx 0.9$ measurement comes from KROSS (Harrison et al., 2017) with $v/\sigma>1$.
@@
However, discrepancies arise when comparing our results to the visual classification results of Ferreira et al. (2023), even after applying similar stellar mass and star formation rate (SFR) cuts. These differences are likely driven by variations in classification schemes and sample selection criteria. In particular, Ferreira et al. studied a broader range of stellar masses, including low-mass systems within $\log(M_{\star}/M_{\odot})\approx[6.6,11.5]$ in the range $6.51.5<z<$ 6.5 for JWST CEERS observations overlapping with the CANDELS EGS field. Our sample is much more massive with $\log(M_{\star}/M_{\odot})=[10,11.4]$. When applying comparable mass cuts, the sample overlap is very limited: we identify only 44 galaxies in the similar low-redshift bin and 75 in the high-redshift bin, with 28 and 41 classified as disks, respectively.
@@
Additionally, differences in classification criteria may contribute to the observed discrepancies. A key distinction in their approach is the “peculiar” category, which includes many galaxies that, in our classification, would likely be labelled as disks with a “clumpy” feature flag or “interacting/merger” or “undefined” when disk classification is uncertain. Another methodological difference lies in the choice of imaging filters: classified galaxies based on images taken in the filter that best matched the rest-frame optical wavelength at the source redshift (i.e., minimizing $\lambda_{\rm rest}/(1+z)$), whereas we utilized imaging across all available filters.
@@ 4.2 Spirals
We use the results from the visual classification of spiral galaxies to assess the spiral fraction in the two redshift bins, as $f_{\mathrm{spiral}}=N_{\mathrm{spiral}}/N_{\mathrm{total}}$. We find spiral fractions of 0.55 and 0.20 in the low- and high-redshift bins, respectively. It is worth noting that the remaining disks in our sample, which are not confidently classified as spirals, could still include galaxies with spiral features. These galaxies might not meet our criteria for confident identification due to factors such as near-edge-on orientations or small angular sizes. Consequently, the inferred spiral fractions could represent only a subset of the real population of disk galaxies with spiral features, as we prioritize robust classification over completeness.
As shown in Figure 10(a), our inferred spiral fractions show general agreement with Kuhn et al. (2024), who also utilize JWST imaging. We compare both their uncorrected and corrected values, where the latter account for observational effects to estimate the intrinsic spiral fraction. While the overall trends are consistent, our results show a steeper decline in the spiral fraction with increasing redshift. As seen in Kuhn et al. (2024), applying corrections for these effects has a larger impact on the inferred intrinsic spiral fractions at higher redshifts. Both our results and those of Kuhn et al. find systematically higher spiral fractions compared to Margalef-Bentabol et al. (2022), whose measurements are based on HST imaging. This discrepancy is likely due to JWST’s superior sensitivity and improved spatial resolution in the near-IR, which enables the identification of more subtle spiral structures that might be missed in HST data. Additionally, JWST’s coverage of longer rest-frame wavelengths enhances the visibility of the underlying disk, boosting confidence in the identification of spirals.
@@
Beyond the overall fraction of spirals, their internal structure also shows striking consistency across redshifts. Applying the same 4/6 voting threshold as in the broad morphological classification, we determined the number of spiral arms in the subset of galaxies that met this criterion. Among the 574 classified galaxies, 343 surpassed this threshold, representing 59.4% of the sample. This includes 260 out of 444 galaxies in the low redshift bin (58.6%) and 81 out of 130 in the high redshift bin (62.3%). We show the number of galaxies as a function of their identified number of spiral arms in Figure 11, with percentages normalized by the number of spiral galaxies in each redshift bin, for which a consensus of 4/6 votes was reached. We find relatively close agreement between the two redshift bins in the normalized arm number fractions, with approximately two-thirds of galaxies exhibiting two arms and about one-third showing three arms in both bins. This suggests little evolution in spiral arm multiplicity across these cosmic epochs.
@@
Figure 11: Distribution of the number of spiral arms identified from visual classification, using a threshold of 4 out of 6 classifiers for agreement. The “Undefined” category refers to systems where classifiers agreed that no specific number of spiral arms could be discerned. The y-axis represents the fraction of spiral galaxies in each redshift bin, normalized by the total number of classified spiral galaxies in that bin. The blue bars correspond to galaxies in the $1.30.8<z<$ 1.3 range, while the orange bars represent galaxies in the $2.52.0<z<$ 2.5 range. Numbers displayed above the bars indicate the absolute counts of galaxies in each category. The green diamonds correspond to the debiased fractions from the luminosity-limited sample of Hart et al. (2016).
@@
Interestingly, our visual classification of spiral arm multiplicity reveals trends broadly consistent with those observed in the local Universe. In both the $1.30.8<z<$ 1.3 and $2.52.0<z<$ 2.5 bins, two-armed spirals remain the dominant morphology, with a fraction comparable to that reported by Hart et al. (2016), who analyzed 62,220 SDSS galaxies from Galaxy Zoo in the redshift range $0.0850.03<z<$ 0.085. For comparison, we use their luminosity-limited subsample ($M_{r}\leq-21$), which spans the same stellar mass range as our data ($\log(M_{\star}/M_{\odot})=[10,11.5]$) and includes debiased fractions to correct for classification biases arising from redshift-dependent effects. Although we do not apply a formal debiasing correction, this adjusted low-redshift baseline provides a useful reference point for interpreting trends at higher redshift.
@@
The agreement in two-arm dominance suggests that grand-design spiral patterns were already common by $z\sim 1$–2 and supports the idea that many high-redshift disks were already cold and massive enough to sustain global $m=2$ modes, either spontaneously, aided by torques introduced by bars or through external drivers such as interactions (e.g., Toomre & Toomre, 1972; Athanassoula, 1984; D’Onghia et al., 2013; Dobbs & Baba, 2014). Although the prevalence of two-armed spirals appears stable, their structural properties may still evolve. For instance, Chugunov et al. (2025) find that pitch angles tend to increase with redshift, suggesting that high-redshift spirals are more loosely wound than their low-redshift counterparts.
@@
Beyond the dominant two-arm population, we find a noticeably higher incidence of three-armed spirals in our sample: while Hart et al. (2016) report that about 20% of local spirals have three arms, we find a higher fraction ($30\%$), pointing to a modest increase in three-fold symmetry at earlier epochs. The relative excess of three-armed systems in our study could be the result of different disk properties at cosmic noon, such as higher gas fractions (e.g., Tacconi et al., 2020). Interestingly, recent simulations have identified the emergence of three-armed spirals in turbulent, gas-rich disks at high redshift, preceding the formation of a two-armed spiral or a bar (Bland-Hawthorn et al., 2024). Notably, one of the earliest high-redshift spirals observed with resolved structure also exhibits a three-armed morphology (Law et al., 2012). In agreement with these results, our results offer the first evidence for an increased incidence of the three-armed spiral morphologies at cosmic noon.
@@ 4.3 Bars
We estimated the bar fractions as $f_{\mathrm{bar}}=N_{\mathrm{bar}}/N_{\mathrm{disk}}$, obtaining $\approx 16\%$ in the low-redshift bin and $\approx 3\%$ in the high-redshift bin. This overall decline with redshift is consistent with theoretical expectations that bars form more readily in dynamically cold, well-settled disks, which become increasingly common at lower redshifts. Supporting this, Sheth et al. (2008) found that bar fractions exceed 60% at $z\approx 0$, but are expected to decline toward earlier epochs due to higher gas fractions and increased turbulence in young disks.
@@
Our measured bar fractions, shown in Figure 10(b), are broadly consistent with those reported in recent high-redshift JWST studies such as Guo et al. (2024) and Le Conte et al. (2024), though we report slightly lower values, particularly in the high-redshift bin. Guo et al. used NIRCam F200W and F [^4] W data from CEERS, combining ellipse fitting and visual inspection to identify bars. Le Conte et al., using CEERS+PRIMER data in F [^4] W, reported higher bar fractions by including both “strongly” and “weakly” barred systems. In their classification scheme, galaxies with at least 3/5 votes for barred were labelled strongly barred, while those with 2/5 votes for barred or at least 3/5 for maybe-barred were considered weakly barred. When considering only the strongly barred subset, their bar fractions drop to 5%–7% for the redshift range $31<z<$ 3, closely matching our measurements within uncertainties.
We also compare our results with Liang et al. (2023), who artificially redshifted nearby galaxies, applying realistic size and luminosity evolution, and simulated NIRCam observations at CEERS depth across multiple filters. Using a bar-size detection threshold of $a_{\mathrm{bar}}>2\times$ the PSF FWHM, their recovered trend in bar fraction also shows a decline with redshift, consistent with our findings.
@@
Despite the overall agreement with previous studies, observational limitations must be carefully considered when interpreting the decline in bar fractions. Surface brightness dimming can obscure low-contrast bars, especially in faint or diffuse galaxies. Lian & Luo (2024) showed that near face-on spirals with faint arms may be misclassified as edge-on systems, further complicating bar identification. In addition, spatial resolution constraints limit the detectability of small-scale bars, particularly at high redshift, where structural features approach the PSF scale and may remain unresolved. In our sample, for instance, bars smaller than 1.5–2 times the FWHM of the F [^4] W PSF are not detected, consistent with the resolution threshold reported by Erwin (2018). These limitations suggest that the observed decline in bar fractions with redshift may reflect not only intrinsic evolution but also increasing challenges in detection at earlier epochs. The true bar fraction at high redshift may be higher than what is currently measurable.
@@
This possibility becomes even more compelling in light of recent observations of rapid radial gas inflows in high-redshift star-forming barred galaxies (Genzel et al., 2023; Arriagada-Neira et al., 2024; Pastras et al. 2025, in prep) and even in dusty star-forming galaxies at $z>3$ (e.g., Umehata et al. 2024; Amvrosiadis et al. 2025. If such inflows are common in early disk galaxies, they could play a crucial role in shaping the evolution of the host galaxies, modifying the conditions for bar instabilities (e.g., Bland-Hawthorn et al., 2024). Moreover, if baryons dominate the local gravitational potential over dark matter in these high-redshift galaxies (e.g., Genzel et al., 2017, 2020; Price et al., 2021; Nestor Shachar et al., 2023), stellar bars could form on rapid timescales of 1–2 Gyr, even in gas-rich turbulent disks (e.g., Fujii et al. 2018; Bland-Hawthorn et al. 2023, 2024), reinforcing the idea that bars at high redshift might be underestimated rather than intrinsically rare.
4.4 Interacting/Mergers

### v2

@@ 4 Morphological evolution
In this section, we present and interpret the results of our visual classification of morphologies and a discussion on the various trends across redshift.
Figure 10: Disk fractions, $f_{\mathrm{disk}}=N_{\mathrm{disk}}/N_{\mathrm{total}}$, as a function of redshift. The dark blue diamonds represent the results from our visual classifications. The light blue diamonds indicate disk fractions when an additional Sérsic index threshold of $n\leq 1.5$ is applied. For comparison, the orange triangles indicate disk fractions from the kinematic analysis of Wisnioski et al. (2015), while the purple squares show measurements from Tiley et al. (2021). The green circles represent disk fractions for the subset of the 3D-HST galaxies in (Ferreira et al., 2023) (F23) that overlap with our sample and satisfy our $M_{\star}$ SFR selection criteria. The gray circles correspond to the unfiltered but still overlapping sample of F23. The gray-shaded regions indicate the redshift bins analyzed in this study.
Figure 11: Spiral fractions defined as $f_{\mathrm{spiral}}=N_{\mathrm{spiral}}/N_{\mathrm{total}}$ for comparison with other studies. Green dots and triangles represent the JWST-based results from (Kuhn et al., 2024), where the white circles show the directly measured spiral fractions, and the filled triangles indicate the values after applying corrections for observational effects. Orange dots represent the fractions inferred from HST imaging in the CANDELS fields by (Margalef-Bentabol et al., 2022).
@@ 4.1 Disks
Figure 10 compares the disk fractions $f_{\mathrm{disk}}=N_{\mathrm{disk}}/N_{\mathrm{total}}$ derived from our study, with the kinematic analysis of KMOS3D (Wisnioski et al., 2015) and a visual morphological classification by Ferreira et al. (2023). We found that our large disk fractions ($0.90\pm 0.03$ and $0.72\pm 0.05$ in the low- and high-redshift bins, respectively) are consistent with the disk fractions estimated from the kinematic analysis of (Wisnioski et al., 2015) across the same redshift range. Additionally, our low-redshift results are also consistent with the findings of Tiley et al. (2021) (KGES) for their sample of 288 galaxies at $1.81.2<z<$ 1.8, where the $z\approx 0.9$ measurement comes from KROSS (Harrison et al., 2017) with $v/\sigma>1$.
@@
However, discrepancies arise when comparing our results to the visual classification results of Ferreira et al. (2023), even after applying similar stellar mass and star formation rate (SFR) cuts. These differences are likely driven by variations in classification schemes and sample selection criteria. In particular, Ferreira et al. studied a broader range of stellar masses, including low-mass systems within $\log(M_{\star}/M_{\odot})\approx[6.6,11.5]$ in the range $6.51.5<z<$ 6.5 for JWST CEERS observations overlapping with the CANDELS EGS field. Our sample is much more massive with $\log(M_{\star}/M_{\odot})=[10,11.4]$. When applying comparable mass cuts, the sample overlap is very limited: we identified only 44 galaxies in the similar low-redshift bin and 75 in the high-redshift bin, where 28 and 41 are classified as disks, respectively.
@@
Additionally, differences in classification criteria may contribute to the observed discrepancies. A key distinction in their approach is the “peculiar” category, which includes many galaxies that, in our classification, would likely be labeled as disks with a clumpy feature flag, interacting/merger, or undefined when disk classification is uncertain. Another methodological difference lies in the choice of imaging filters: classified galaxies based on images taken in the filter that best matched the rest-frame optical wavelength at the source redshift (i.e., minimizing $\lambda_{\rm rest}/(1+z)$), whereas we utilized imaging across all available filters.
In addition to disk fractions, the classifications of galaxy morphologies also allow the measurement of the fraction of disks that appear featureless, i.e., lacking bars, bulges, clumps, or spiral arms. A recent study by Smethurst et al. (2025) used Galaxy Zoo (Lintott et al., 2008) classifications to analyze this population at $z\approx 3-7$, finding that although many galaxies are disk-like, most lack visible internal structure. To enable a direct comparison with our sample, we focused on their reported featureless disk fractions in the redshift range $z\approx 0.8-2.5$, which span fractions of $\approx 0.75-0.59$ with a mild decline toward lower redshift. After estimating a correction for observational constraints and redshift effects, they infer that the intrinsic featureless disk fractions are more likely in the range $\approx 0.38-0.29$. In our visual classifications, based on six expert classifiers and a different morphological scheme, we found featureless [^11] Featureless refers to galaxies that were not classified as having bulges, clumps, bars, or spiral arms, to approximate the definition used by Smethurst et al., despite differences in classification schemes and criteria. disk fractions of 0.47 and 0.20 in the high- and low-redshift bins, respectively. This redshift trend, opposite to that of Smethurst et al. (2025), may reflect differences in classification methodology, image depth, or sample selection: our study targets more massive galaxies at cosmic noon, whereas theirs includes a broader population over a wider redshift range. Nonetheless, the average values are remarkably similar (0.34 in our case vs. 0.32 from their bias-corrected estimates), suggesting that despite methodological differences, the underlying physical picture may be broadly consistent.
@@ 4.2 Spirals
We use the results from the visual classification of spiral galaxies to assess the spiral fraction in the two redshift bins, as $f_{\mathrm{spiral}}=N_{\mathrm{spiral}}/N_{\mathrm{total}}$. We found spiral fractions of 0.55 and 0.20 in the low- and high-redshift bins, respectively. It is worth noting that the remaining disks in our sample, which are not confidently classified as spirals, could still include galaxies with spiral features. These galaxies might not meet our criteria for confident identification due to factors such as near-edge-on orientations or small angular sizes. Consequently, the inferred spiral fractions could represent only a subset of the real population of disk galaxies with spiral features, as we prioritized robust classification over completeness.
As shown in Figure 11, our inferred spiral fractions show general agreement with Kuhn et al. (2024), who also utilize JWST imaging. We compare both their uncorrected and corrected values, where the latter account for observational effects, to estimate the intrinsic spiral fraction. While the overall trends are consistent, our results show a steeper decline in the spiral fraction with increasing redshift. As seen in Kuhn et al. (2024), applying corrections for these effects has a larger impact on the inferred intrinsic spiral fractions at higher redshifts. Both our results and those of Kuhn et al. find systematically higher spiral fractions compared to Margalef-Bentabol et al. (2022), whose measurements are based on HST imaging. This discrepancy is likely due to JWST’s superior sensitivity and improved spatial resolution in the near-IR, which enables the identification of more subtle spiral structures that might be missed in HST data. Additionally, JWST’s coverage of longer rest-frame wavelengths enhances the visibility of the underlying disk, boosting confidence in the identification of spirals.
@@
Beyond the overall fraction of spirals, their internal structure also shows striking consistency across redshifts. Applying the same 4/6 voting threshold as in the broad morphological classification, we determined the number of spiral arms in the subset of galaxies that met this criterion. Among the 574 classified galaxies, 341 surpassed this threshold, representing 60% of the sample. This includes 260 out of 444 galaxies in the low-redshift bin (59%) and 81 out of 130 in the high-redshift bin (62%). We show the number of galaxies as a function of their identified number of spiral arms in Figure 12, where the percentages are normalized by the number of spiral galaxies in each redshift bin, for which a consensus of 4/6 votes was reached. We found relatively close agreement between the two redshift bins in the normalized arm number fractions: approximately two-thirds of galaxies exhibit two arms and about one-third show three arms in both bins. This suggests little evolution in spiral arm multiplicity across these cosmic epochs.
@@
Figure 12: Distribution of the number of spiral arms identified from visual classification, using a threshold of four out of six classifiers for agreement. The undefined category refers to systems where classifiers agreed that no specific number of spiral arms could be discerned. The y-axis represents the fraction of spiral galaxies in each redshift bin, normalized by the total number of classified spiral galaxies in that bin. The blue bars correspond to galaxies in the $1.30.8<z<$ 1.3 range, while the orange bars represent galaxies in the $2.52.0<z<$ 2.5 range. Numbers displayed above the bars indicate the absolute counts of galaxies in each category. The green diamonds correspond to the debiased fractions from the luminosity-limited sample of Hart et al. (2016).
@@
Interestingly, our visual classification of spiral arm multiplicity reveals trends broadly consistent with those observed in the local Universe. In both the $1.30.8<z<$ 1.3 and $2.52.0<z<$ 2.5 bins, two-armed spirals remain the dominant morphology, and have fractions comparable to that reported by Hart et al. (2016), who analyzed 62,220 Sloan Digital Sky Survey (SDSS; York et al., 2000) galaxies as part of the Galaxy Zoo project in the redshift range $0.0850.03<z<$ 0.085. For comparison, we used their luminosity-limited subsample ($M_{r}\leq-21$), which spans the same stellar mass range as our data ($\log(M_{\star}/M_{\odot})=[10,11.5]$) and includes debiased fractions to correct for classification biases arising from redshift-dependent effects. Although we did not apply a formal debiasing correction, this adjusted low-redshift baseline provides a useful reference point for interpreting trends at higher redshift.
@@
The agreement in two-arm dominance suggests that grand-design spiral patterns were already common by $z\sim 1$–2 and supports the idea that many high-redshift disks were already cold and massive enough to sustain global $m=2$ modes, either spontaneously or aided by torques introduced by bars, or through external drivers such as interactions (e.g., Toomre & Toomre, 1972; Athanassoula, 1984; D’Onghia et al., 2013; Dobbs & Baba, 2014). Although the prevalence of two-armed spirals appears stable, their structural properties may still evolve. For instance, Chugunov et al. (2025) found that pitch angles tend to increase with redshift, suggesting that high-redshift spirals are more loosely wound than their low-redshift counterparts.
@@
Beyond the dominant two-arm population, we found a noticeably higher incidence of three-armed spirals in our sample: while Hart et al. (2016) reported that a fraction of about 0.20 of local spirals have three arms, we found a higher fraction ($\approx 0.30$), pointing to a modest increase in three-fold symmetry at earlier epochs. The relative excess of three-armed systems in our study could be the result of different disk properties at cosmic noon, such as higher gas fractions (e.g., Tacconi et al., 2020). Interestingly, recent simulations have identified the emergence of three-armed spirals in turbulent, gas-rich disks at high redshift, preceding the formation of a two-armed spiral or a bar (Bland-Hawthorn et al., 2024). Notably, one of the earliest high-redshift spirals observed with resolved structure also exhibits a three-armed morphology (Law et al., 2012). In agreement with these results, our results offer the first evidence for an increased incidence of the three-armed spiral morphologies at cosmic noon.
@@ 4.3 Bars
Figure 13: Bar fractions ($f_{\mathrm{bar}}=N_{\mathrm{bar}}/N_{\mathrm{disk}}$) from the subsample of disks with $e<0.5$ where the isophotal fitting analysis is reliable. The orange dots show the results from Guo et al. (2025), while the green dots correspond to the fractions inferred by Le Conte et al. (2024), where the green triangles indicate the fractions inferred using only their sample of strongly barred galaxies. The purple dots indicate the Galaxy-Zoo based measurements from (Géron et al., 2025) using JWST-NIRCam imaging. The gray triangles represent simulated bar fractions from Liang et al. (2024) for a set of JWST filters, given an intrinsic $f_{\mathrm{bar}}$ of 0.68 at $z\approx 0$. Finally, the red and gray dots indicate the HST-based bar fractions from (Simmons et al., 2014) and (Melvin et al., 2014), respectively.
@@
We estimated the bar fractions as $f_{\mathrm{bar}}=N_{\mathrm{bar}}/N_{\mathrm{disk}}$, obtaining $\approx 0.16\pm 0.05$ in the low-redshift bin and $\approx 0.03\pm 0.03$ in the high-redshift bin. This overall decline with redshift is consistent with theoretical expectations that bars form more readily in dynamically cold, well-settled disks, which become increasingly common at lower redshifts. Supporting this, Sheth et al. (2008) found that bar fractions exceed 0.60 at $z\approx 0$, but are expected to decline toward earlier epochs due to higher gas fractions and increased turbulence in young disks. Stellar population studies of local barred galaxies also support this physical picture: de Sá-Freitas et al. (2025) inferred that many bars were already in place at high redshift, showing cumulative bar fractions at a given redshift that are broadly consistent with morphological studies, based on their backward extrapolations.
Further context comes from Galaxy Zoo classifications based on both HST and JWST imaging. Early studies using HST data, including Melvin et al. (2014) and Simmons et al. (2014) reported bar fractions that already showed a decline in $f_{\mathrm{bar}}$ as a function of redshift up to $z\approx 1.7$. More recently, Géron et al. (2025) extended Galaxy Zoo classifications to JWST NIRCam imaging and found systematically higher bar fractions compared to the HST-based studies at similar redshifts but a similar decreasing trend. The step up from HST to JWST , while still consistent with a declining trend, suggests that part of the observed evolution is driven by detection biases. The improved resolution and sensitivity of JWST allow the identification of fainter and smaller bars that were missed in earlier classifications. As shown in Figure 13, our results lie between the HST and JWST Galaxy Zoo measurements, closely matching the latter in the low-redshift bin, but falling below it in the high-redshift regime. While our visual classification approach differs from the Galaxy Zoo citizen science framework in both scope and methodology, the agreement at low redshift and divergence at higher redshift further highlight the influence of detection thresholds and classification systematics on $f_{\mathrm{bar}}$ estimates, effects that become increasingly significant at high redshift.
@@
Our measured bar fractions align more closely with other JWST studies that use expert classifications or ellipse fitting, such as Guo et al. (2025) and Le Conte et al. (2024). This likely reflects methodological similarities, in contrast to the Galaxy Zoo approach used by Géron et al. (2025), which adopts a different classification framework. Guo et al. used NIRCam F200W and F [^4] W data from CEERS, combining ellipse fitting and visual inspection to identify bars. Le Conte et al., using CEERS+PRIMER data in F [^4] W, reported higher bar fractions by including both strongly and weakly barred systems. In their classification scheme, galaxies with at least 3/5 votes for barred were labeled strongly barred, while those with 2/5 votes for barred or at least 3/5 for maybe-barred were considered weakly barred. When considering only the strongly barred subset, their bar fractions drop to 0.05–0.07 for the redshift range $31<z<$ 3, closely matching our measurements within the uncertainties.
Even with the improved resolution and sensitivity of JWST, bar detectability remains strongly dependent on the choice of filter and its corresponding rest-frame wavelength coverage. Liang et al. (2024) demonstrated this by artificially redshifting local galaxies and simulating JWST/NIRCam observations across multiple filters, incorporating realistic size and luminosity evolution. Adopting a conservative detection threshold of $a_{\mathrm{bar}}>2\times$ the PSF FWHM, they recovered a declining trend in bar fraction with redshift, but also found systematic offsets as a function of filter, reporting higher bar fractions in bluer bands and lower values in redder bands, as illustrated in Figure 13.
@@
Additional observational limitations must be carefully considered when interpreting the decline in bar fractions. Surface brightness dimming can obscure low-contrast bars, especially in faint or diffuse galaxies. Lian & Luo (2024) showed that near face-on spirals that have faint arms may be misclassified as edge-on systems, further complicating bar identification. In addition, spatial resolution constraints limit the detectability of small-scale bars, particularly at high redshift, where structural features approach the PSF scale and may remain unresolved. In our sample, for instance, bars smaller than 1.5–2 times the FWHM of the F [^4] W PSF are not detected, consistent with the resolution threshold reported by Erwin (2018). These limitations suggest that the observed decline in bar fractions with redshift may reflect not only intrinsic evolution but also increasing challenges in detection at earlier epochs. The true bar fraction at high redshift may be higher than what is currently measurable.
@@
This possibility becomes even more compelling in light of recent observations of rapid radial gas inflows in high-redshift star-forming barred galaxies (Genzel et al., 2023; Arriagada-Neira et al., 2025; Pastras et al., 2025) and even in dusty star-forming galaxies at $z>3$ (e.g., Umehata et al. 2025; Amvrosiadis et al. 2025. If such inflows are common in early disk galaxies, they could play a crucial role in shaping the evolution of the host galaxies, modifying the conditions for bar instabilities (e.g., Bland-Hawthorn et al., 2024), potentially leading to an increase in the intrinsic bar fractions. Moreover, if baryons dominate the local gravitational potential over dark matter in these high-redshift galaxies (e.g., Genzel et al., 2017, 2020; Price et al., 2021; Nestor Shachar et al., 2023), stellar bars could form on rapid timescales of 1–2 Gyr, even in gas-rich turbulent disks (e.g., Fujii et al. 2018; Bland-Hawthorn et al. 2023, 2024).
Complementary insights come from cosmological simulations, which generally predict a relatively constant intrinsic bar fraction across cosmic time. For example, Rosas-Guevara et al. (2022), using the TNG50 simulation suite (Pillepich et al., 2019), find that bar fractions remain roughly flat at $\sim 0.40$ from $z=0$ to $z\approx 2$. Similarly, earlier work by Zhao et al. (2020), based on TNG100 (Nelson et al., 2019), found an intrinsically constant, but observationally dropping bar fraction up to $z\sim 1$. Independently, Frosst et al. (in prep.), using the COLIBRE cosmological hydrodynamical simulations (Schaye et al., in prep.), report a similarly flat bar fraction over this redshift range. Taken together, cosmological simulations and emerging observational evidence for bar-driven inflows at high redshift support the view that the observed decline in bar fractions is partly real but significantly amplified by detection limitations, leaving many high-redshift bars potentially hidden.
4.4 Interacting and merger systems